TheSmartOne (thesmartone):

How would I find this limit analytically? Any hints would be appreciated. :)

TheSmartOne (thesmartone):

\[\lim_{x \rightarrow \frac{\pi}{4}}\left(\frac{1-\tan~\theta}{\sin~\theta-\cos~\theta}\right)\]

OpenStudy (tay1lynn):

umm,,,

OpenStudy (tay1lynn):

@agent0smith

TheSmartOne (thesmartone):

If I divide everything in the denominator by cos\(\theta\) I get tan\(\theta\) - 1 But I don't know if that would help any, mhmm

OpenStudy (holsteremission):

What do you mean by "analytically"? As in no L'Hopital's rule or series expansions allowed?

OpenStudy (irishboy123):

\(\Large \lim_{\theta \rightarrow \frac{\pi}{4}}\left(\frac{1-\tan~\theta}{\sin~\theta-\cos~\theta}\right)\) \(= \Large \lim_{\theta \rightarrow \frac{\pi}{4}} \sec \theta \left(\frac{\cos \theta -\sin~\theta}{\sin~\theta-\cos~\theta}\right)\)

TheSmartOne (thesmartone):

I haven't learned that yet. Analytically here would be factoring the expression so that way we could get rid of the restriction of substituting pi/4 in and finding the limit, mhmm

TheSmartOne (thesmartone):

Oops, it was a typo. x was supposed to be theta

OpenStudy (irishboy123):

done?

OpenStudy (zzr0ck3r):

I am not sure what you mean. I see the term analytically as using analysis. You don't really use analysis to find the limit but to prove the limit. i.e. A sequence has a limit \(a\) iff \(\forall \epsilon \mathbb{R} \ \ \exists N\in \mathbb{N} \ \ \forall n\in \mathbb{N} \ \ |x_n-a|<\epsilon\)

OpenStudy (irishboy123):

me too, mate

TheSmartOne (thesmartone):

There are other questions that ask to prove the limit using the epsilon-delta defintion of the limit and they say so but this question asks to find the limit analytically

TheSmartOne (thesmartone):

\[\lim_{x \rightarrow \frac{\pi}{4}}\left(\frac{1-\tan~\theta}{\sin~\theta-\cos~\theta}\right)\] \[\lim_{x \rightarrow \frac{\pi}{4}}\left(\sec \theta \cdot \frac{\cos\theta-\sin \theta}{\sin~\theta-\cos~\theta}\right)\] \[\lim_{x \rightarrow \frac{\pi}{4}}\left(\sec \theta \cdot \frac{\cos\theta-\sin \theta}{-\cos \theta + \sin \theta }\right)\] \[\lim_{x \rightarrow \frac{\pi}{4}}\left(\sec \theta \cdot -1\right)\] \[\lim_{x \rightarrow \frac{\pi}{4}}\left(-\sec \theta \right)\]

jhonyy9 (jhonyy9):

TSO sorry that will be a bad idea but what may be result than rewrite (cos theta -sin theta)/(sin theta -cos theta) = - (sin theta -cos theta)/(sin theta -cos theta) = - 1

TheSmartOne (thesmartone):

Yes, thank you Jhonny. It took me a while to realize that :)

jhonyy9 (jhonyy9):

was my pleasure so this may be usefully ?

TheSmartOne (thesmartone):

so sec theta = 1/cos theta cos (pi/4) = 1/sqrt(2) sec(pi/4) = sqrt(2) -sec(pi/4) = -sqrt(2) soooo \[\lim_{x \rightarrow \frac{\pi}{4}}\left(\frac{1-\tan~\theta}{\sin~\theta-\cos~\theta}\right) =\boxed{\bf{ -\sqrt{2}}}\]

TheSmartOne (thesmartone):

Thanks everyone :)